# Underwater Camouflaged Object Tracking Meets Vision-Language SAM2

### v3

@@ Abstract
Over the past decade, significant progress has been made in visual object tracking, largely due to the availability of large-scale datasets. However, these datasets have primarily focused on open-air scenarios and have largely overlooked underwater animal tracking—especially the complex challenges posed by camouflaged marine animals. To bridge this gap, we take a step forward by proposing the first large-scale multi-modal underwater camouflaged object tracking dataset, namely UW-COT220. Based on the proposed dataset, this work first comprehensively evaluates current advanced visual object tracking methods, including SAM- and SAM2-based trackers, in challenging underwater environments, e.g., coral reefs. Our findings highlight the improvements of SAM2 over SAM, demonstrating its enhanced ability to handle the complexities of underwater camouflaged objects. Furthermore, we propose a novel vision-language tracking framework called VL-SAM2, based on the video foundation model SAM2. Experimental results demonstrate that our VL-SAM2 achieves state-of-the-art performance on the UW-COT220 dataset. The dataset and codes are available at here.
@@ 1 Introduction
Visual object tracking (VOT) involves continuously locating a target object within a video sequence and has applications in animal monitoring and conservation [34, 36], autonomous vehicles [31, 25], and surveillance [39, 8], etc. Its significance lies in enabling machines to perceive and interpret dynamic environments, supporting tasks like animal behavioral analysis [38] and decision-making [22]. Although substantial progress has been achieved in terrestrial and open-air scenarios [16, 9], tracking in underwater environments remains highly challenging, largely limiting the effectiveness of conventional algorithms [38, 2, 1]. Consequently, there is a pressing need for specialized datasets and methods to tackle the complexities of underwater tracking, especially when objects blend with their surroundings, known as camouflaged objects [4, 6]. However, despite its importance, underwater camouflaged object tracking remains an unexplored field.
Underwater object tracking presents numerous unique challenges due to light attenuation [38, 2], scattering effects [1, 22], and frequent camouflage scenarios [6, 4], which significantly degrade the performance of conventional tracking paradigms. Correlation filter-based methods [32, 12, 11, 33], while computationally efficient, fail to handle the dynamic appearance variations caused by underwater optical distortions. Siamese networks [3, 20] struggle with feature representation learning in low-contrast and turbid environments. Transformer-based approaches [31, 5, 28], though powerful in modeling long-range dependencies, often suffer from attention distraction due to cluttered underwater backgrounds. Even the emerging Mamba-based methods [15, 19, 40], despite their efficient sequence modeling, lack the specialized capability to address underwater-specific degradations. In contrast, SAM- and SAM2-based trackers [35, 37, 10, 30, 29] demonstrate superior robustness by leveraging their pre-trained visual foundation model’s strong generalization ability and prompt-guided segmentation mechanism. This enables effective handling of camouflaged marine animals through semantic-aware feature extraction and adaptive multi-scale processing, making SAM- and SAM2-based approaches particularly suitable for challenging underwater tracking scenarios.
@@
Recently, several video camouflaged object detection and segmentation datasets have been proposed [4, 6, 14], but they primarily focus on open-air scenarios. Following the latest large-scale underwater object tracking dataset, i.e., WebUOT-1M [38], we take a step forward and construct the first multi-modal underwater camouflaged object tracking dataset, UW-COT220 (see Fig. 1 and Tab. 1), as a benchmark for evaluating and developing advanced underwater camouflaged object tracking methods. The dataset consists of 220 underwater video sequences, spanning 96 categories, with approximately 159 K frames. We annotate each video with a language description to promote the exploration of multi-modal underwater camouflaged object tracking. As a single-object visual tracking task, we provide bounding box annotations for the camouflaged underwater objects in each frame. Additionally, using manually annotated bounding boxes as prompts, we generate mask annotations for the camouflaged underwater objects. To advance research in the related field, we propose a simple yet effective VL tracking framework, called VL-SAM2. By releasing our dataset and codes, we aim to inspire the exploration of diverse underwater vision tasks and foster the development of underwater camouflaged object tracking and beyond.
Our key contributions are as follows: 1) We introduce UW-COT220, the first large-scale multi-modal underwater camouflaged object tracking dataset with bounding boxes, masks, and language descriptions. 2) We propose VL-SAM2, a new VL tracker based on SAM2, leveraging language prompts and motion-aware prediction for robust underwater tracking. 3) Our VL-SAM2 outperforms state-of-the-art (SOTA) tracking methods, proving its effectiveness in handling underwater camouflage and complex scenes.
@@ 2 UW-COT220 Dataset
Our goal is to construct a large-scale underwater camouflaged object tracking dataset that involves a wide variety of categories and various real underwater scenes for evaluating and developing general underwater camouflaged object tracking methods. To achieve this, we collect underwater videos from video-sharing platforms (e.g., YouTube [^1] https://www.youtube.com/) and existing tracking datasets (e.g., WebUOT-1M [38] and VastTrack [22]), and we filter out 220 videos that contain underwater camouflaged objects. Following existing datasets [34, 22, 38], we provide a bounding box annotation for the camouflaged object in each frame, represented as $[x,y,w,h]$, where $x$ and $y$ denote the coordinates of the top-left corner of the object, and $w$ and $h$ indicate the width and height of the object. In each video, we annotate an English sentence describing the color, behavior, attributes, and surroundings of the target. Furthermore, to enhance the accuracy of object representation, we use the SAM-powered interactive semi-automatic annotation tool [17] to generate mask annotations for the camouflaged objects. Some examples and detailed statistics of the UW-COT220 dataset are provided in Fig. 1 and Tab. 1. To the best of our knowledge, UW-COT220 is the first large-scale multi-modal benchmark for underwater camouflaged object tracking, featuring a diverse set of 96 categories and various challenging underwater scenes.
@@ 3 Proposed Method
In this section, we introduce vision-language (VL)-SAM2, a new vision-language tracking framework for multi-modal underwater camouflaged object tracking. VL-SAM2 (see Fig. 2) is based on the latest video foundation model, SAM2 [29], and extends its capabilities to underwater scenarios. VL-SAM2 consists of three main components: a visual branch (i.e., SAM2), a language branch (LB), and a motion-aware target prediction (MATP) module [38].
We use Hiera-L [26] as the image encoder and CLIP ViT-B/32 [23] as the language encoder. During training, we freeze the parameters of both the image encoder and the language encoder, and fine-tune the other modules. This allows us to train the proposed VL-SAM2 on an Ubuntu server with 8 A6000 GPUs. Our model was trained on the Ref-YouTube-VOS [27] dataset and evaluated in a zero-shot manner on the UW-COT220 dataset.
@@ 5 Conclusion
In this work, we introduce UW-COT220, the first large-scale multi-modal underwater camouflaged object tracking dataset. To promote related research, we propose a new vision-language tracking framework VL-SAM2. Future work includes expanding the scale and modalities of our dataset and exploring more underwater vision tasks.

### v5

@@ Abstract
Over the past decade, significant progress has been made in visual object tracking, largely due to the availability of large-scale datasets. However, these datasets have primarily focused on open-air scenarios and have largely overlooked underwater animal tracking—especially the complex challenges posed by camouflaged marine animals. To bridge this gap, we take a step forward by proposing the first large-scale multi-modal underwater camouflaged object tracking dataset, namely UW-COT220. Based on the proposed dataset, this work first comprehensively evaluates current advanced visual object tracking methods, including SAM- and SAM2-based trackers, in challenging underwater environments, e.g., coral reefs. Our findings highlight the improvements of SAM2 over SAM, demonstrating its enhanced ability to handle the complexities of underwater camouflaged objects. Furthermore, we propose a novel vision-language tracking framework called VL-SAM2, based on the video foundation model SAM2. Extensive experimental results demonstrate that the proposed VL-SAM2 achieves state-of-the-art performance across underwater and open-air object tracking datasets. The dataset and codes are available at https://github.com/983632847/Awesome-Multimodal-Object-Tracking.
@@ 1 Introduction
Visual object tracking (VOT) involves continuously locating a target object within a video sequence and has applications in animal monitoring and conservation [21, 43, 45], autonomous vehicles [40, 29], and surveillance [48, 10], etc. Its significance lies in enabling machines to perceive and interpret dynamic environments, supporting tasks like animal behavioral analysis [33, 6, 47] and decision-making [26, 18]. Although substantial progress has been achieved in terrestrial and open-air scenarios [11, 18, 10], tracking in underwater environments remains highly challenging, largely limiting the effectiveness of conventional algorithms [47, 2, 1]. Consequently, there is a pressing need for specialized datasets and methods to tackle the complexities of underwater tracking, especially when objects blend with their surroundings, known as camouflaged objects [4, 7]. However, despite its importance, underwater camouflaged object tracking remains an unexplored field.
Underwater object tracking presents numerous unique challenges due to light attenuation [47, 2], scattering effects [1, 26], and frequent camouflage scenarios [7, 4], which significantly degrade the performance of conventional tracking paradigms. Correlation filter-based methods [37, 41, 14, 13, 42], while computationally efficient, fail to handle the dynamic appearance variations caused by underwater optical distortions. Siamese networks [9, 3, 23, 15] struggle with feature representation learning in low-contrast and turbid environments. Transformer-based approaches [25, 36, 34, 40, 5, 35, 49, 48], though powerful in modeling long-range dependencies, often suffer from attention distraction due to cluttered underwater backgrounds. Even the emerging Mamba-based methods [17, 22, 50], despite their efficient sequence modeling, lack the specialized capability to address underwater-specific degradations. In contrast, SAM- and SAM2-based trackers [44, 46, 12, 39, 38, 32] demonstrate superior robustness by leveraging their pre-trained visual foundation model’s strong generalization ability and prompt-guided segmentation mechanism. This enables effective handling of camouflaged marine animals through semantic-aware feature extraction and adaptive multi-scale processing, making SAM- and SAM2-based approaches particularly suitable for challenging underwater tracking scenarios.
@@
Recently, several video camouflaged object detection and segmentation datasets have been proposed [4, 7, 16], but they primarily focus on open-air scenarios. Following the latest large-scale underwater object tracking dataset, i.e., WebUOT-1M [47], we take a step forward and construct the first multi-modal underwater camouflaged object tracking dataset, UW-COT220 (see Fig. 1 and Tab. 1), as a benchmark for evaluating and developing advanced underwater camouflaged object tracking methods. The dataset consists of 220 underwater video sequences, spanning 96 categories, with approximately 159 K frames. We annotate each video with a language description to promote the exploration of multi-modal underwater camouflaged object tracking. As a single-object visual tracking task, we provide bounding box annotations for the camouflaged underwater objects in each frame. To support precise tracking, we annotate each frame with absent labels: “0” for target presence and “1” for full occlusion or out-of-view. Additionally, using manually annotated bounding boxes as prompts, we generate mask annotations for the camouflaged underwater objects. To ensure the quality of the annotations, we have carefully checked the bounding boxes, masks, and language descriptions for multiple rounds. To advance research in the related field, we propose a simple yet effective VL tracking framework, called VL-SAM2. By releasing our dataset and codes, we aim to inspire the exploration of diverse underwater vision tasks and foster the development of underwater camouflaged object tracking and beyond.
Our main contributions can be summarized as follows:
  • We introduce UW-COT220, the first large-scale multi-modal underwater camouflaged object tracking dataset with bounding boxes, masks, and language descriptions.
  • We propose VL-SAM2, a new VL tracker based on SAM2, leveraging language prompts and motion-aware prediction for robust underwater tracking.
  • Our VL-SAM2 outperforms state-of-the-art (SOTA) tracking methods, proving its effectiveness in handling underwater camouflage and complex scenes.
@@ 2 UW-COT220 Dataset
Our goal is to construct a large-scale underwater camouflaged object tracking dataset that involves a wide variety of categories and various real underwater scenes for evaluating and developing general underwater camouflaged object tracking methods. To achieve this, we collect underwater videos from video-sharing platforms (e.g., YouTube [^1] https://www.youtube.com/) and existing tracking datasets (e.g., WebUOT-1M [47] and VastTrack [26]), and we filter out 220 videos that contain underwater camouflaged objects. Note that WebUOT-1M [47] is currently the largest general-purpose underwater object tracking dataset, comprising 81 underwater camouflage videos. We further construct a large-scale dataset dedicated to underwater camouflaged object tracking, featuring more diverse underwater scenes and comprehensive annotations. Following existing datasets [43, 26, 47], we provide a bounding box annotation for the camouflaged object in each frame, represented as $[x,y,w,h]$, where $x$ and $y$ denote the coordinates of the top-left corner of the object, and $w$ and $h$ indicate the width and height of the object. In each video, we annotate an English sentence describing the color, behavior, attributes, and surroundings of the target. Furthermore, to enhance the accuracy of object representation, we use the SAM-powered interactive semi-automatic annotation tool [19] to generate mask annotations for the camouflaged objects. Some examples and detailed statistics of the UW-COT220 dataset are provided in Fig. 1 and Tab. 1. To the best of our knowledge, UW-COT220 is the first large-scale multi-modal benchmark for underwater camouflaged object tracking, featuring a diverse set of 96 categories and various challenging underwater scenes.
@@ 3 Proposed Method
In this section, we introduce vision-language (VL)-SAM2, a new vision-language tracking framework for multi-modal underwater camouflaged object tracking. The proposed VL-SAM2 (see Fig. 2) is based on the latest video foundation model, SAM2 [38], and extends its capabilities to underwater scenarios. VL-SAM2 consists of three main components, including a visual branch (i.e., SAM2), a language branch (LB), and a motion-aware target prediction (MATP) module [47].
We use Hiera-L [30] as the image encoder and CLIP ViT-B/32 [27] as the language encoder. During training, we freeze the parameters of both the image encoder and the language encoder, and fine-tune the other modules. This allows us to train the proposed VL-SAM2 on an Ubuntu server with 8 A6000 GPUs. Our model was trained on the Ref-YouTube-VOS [31] dataset and evaluated in a zero-shot manner on the UW-COT220 and existing datasets [10, 47].
@@ 5 Conclusion
In this work, we introduce UW-COT220, the first large-scale multi-modal underwater camouflaged object tracking dataset. To promote related research, we propose a new vision-language tracking framework VL-SAM2. Experimental results show that our proposed VL-SAM2 exhibits excellent zero-shot generalization capability in both underwater and open-air scenarios. Future work includes expanding the scale and modalities of our dataset and exploring more underwater vision tasks.
Acknowledgements. This work was supported by the National Natural Science Foundation of China (No. 62471420), GuangDong Basic and Applied Basic Research Foundation (2025A1515012296), CCF-Tencent Rhino-Bird Open Research Fund, and the Major Project of Technology Innovation and Application Development of Chongqing (CSTB2023TIAD-STX0015).